A hydraulic lift has pistons with diameters 8.00 cm and 36.0 cm, respectively. If a force of 825 N is exerted at the input piston, what maximum mass can be lifted by the output piston


1
Expert's answer
2022-01-30T13:51:18-0500

Given:

"d_1=8.00\rm\: cm"

"d_2=36.0\rm\: cm"

"F_1=825\:\rm N"


The hydraulic press equation says

"\frac{F_1}{A_1}=\frac{F_2}{A_2}"

Hence, the output force

"F_2=F_1\frac{A_2}{A_1}=F_1\frac{d_2^2}{d_1^2}"

The maximum mass can be lifted by the output piston

"m=\frac{F_2}{g}=\frac{F_1}{g}\frac{d_2^2}{d_1^2}"

"m=\frac{825}{9.81}\frac{36.0^2}{8.00^2}=1700\:\rm kg"